OpenStudy (anonymous):

Input in standard form the equation of the given line. The line through (0, -3) and (3, 0)

OpenStudy (anonymous):

say given points are A(x1,y1) and B(x2, y2). then the equation of line passing through A and B is (y-y1)/(y2-y1) = (x-x1)/(x2-x1)

OpenStudy (saifoo.khan):

try otherwise i will help!

OpenStudy (anonymous):

help:(

OpenStudy (saifoo.khan):

lol, alright. find the slope: 0-(-3) ----- 3 - 0 equals 1

OpenStudy (saifoo.khan):

y = mx + c -3 = 1(0) + c c = -3

OpenStudy (saifoo.khan):

now, y = mx + c y = x -3

OpenStudy (anonymous):

given points are (x1,y1) = (0,-3) and (x2, y2) = (3,0). So x1=0, y1=-3, x2=3, y2=0; that gives (y-(-3))/(0-(-3)) = (x-0)/(3-0) (y+3)/3 = x/3 y+3=x y=x-3